What is the oxidation state of Cl in each of the following ions?


Part A ClO-


Part B ClO2-   


Part C ClO3-


Part D ClO4-
Answer #1

A= 1, b= 3, c= 5, d= 7
Answer #2

In oxyanions of Chlorine, the oxidation state of Oxygen atom is -2. Let the oxidation number of Cl in each of the anions be X.

The sum of the oxidation states of the different atoms in the compound should be equal to the total charge on the anion.

1. X + (-2) = -1

X - 2 = -1

X = +1


2. X + 2*(-2) = -1

X - 4 = -1

X = +3


3. X + 3*(-2) = -1

X - 6 = -1

X = +5


4. X + 4*(-2) = -1

X - 8 = -1

X = +7
Answer #3

+1


0

-1

-4